OpenStudy (anonymous):

What are the solutions of 3x^2 + 6x + 6 = 0?

OpenStudy (anonymous):

start by dividing by 3, get \[x^2+2x+2=0\] you will get two complex solutions

OpenStudy (anonymous):

@satellite73 ok?

OpenStudy (anonymous):

if you are allowed complex solutions you can write \[x^2+2x=-2\] \[(x+1)^2=-2+1=-1\] \[x+1=\pm i\] \[x=-1\pm i\]

OpenStudy (anonymous):

@satellite73 my choices are, –2 ± i –2 ± 2i –1 ± i 1 ± i I can only pick one it's the last one right?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

as usual it is C

OpenStudy (anonymous):

@satellite73 as usual? lol

OpenStudy (anonymous):

@satellite73 can you help me with another question that I have?